# Eaman Jahani

### v5

@@ G.1 Hypotheses and Results
  • CLIP embedding cosine similarity
    • Z-score rescaling: We found statistically significant evidence for performance differences across treatment arms when examining participants’ first attempts, average performance across all attempts, and best attempts (all measured using taskwide Z-scores rather than task-iteration Z-scores).
      For first attempt performance, DALL-E 3 (Verbatim) participants significantly outperformed both DALL-E 3 (Revised) and DALL-E 2 participants, while no statistically significant difference emerged between DALL-E 3 (Revised) and DALL-E 2 participants.
      Average performance across all attempts showed the same pattern: DALL-E 3 (Verbatim) outperformed both other conditions, with no significant difference between DALL-E 3 (Revised) and DALL-E 2.
@@
      When comparing participants’ best attempts, we found statistically significant differences between all three treatment arms in a clear hierarchy: DALL-E 3 (Verbatim) produced better results than DALL-E 3 (Revised), which in turn produced better results than DALL-E 2. ANOVA tests for all three performance measures (first, average, and best attempts) confirmed significant effects of model assignment.
@@
    • Percentile rank rescaling: Results using percentile rank rescaling mirrored those found with Z-score rescaling. DALL-E 3 (Verbatim) participants outperformed both other conditions on first attempt performance and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2 for these metrics. For best attempt performance, all pairwise comparisons revealed significant differences, with DALL-E 3 (Verbatim) outperforming DALL-E 3 (Revised), which outperformed DALL-E 2. ANOVA tests confirmed significant effects of model assignment for all three performance measures.
  • DreamSim
@@
    • Z-score rescaling: Results using DreamSim with Z-score rescaling showed the same pattern as CLIP embedding cosine similarity. DALL-E 3 (Verbatim) participants significantly outperformed both other conditions on first attempt and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2. For best attempts, we again found the same hierarchical pattern of performance differences across all three conditions. ANOVA tests confirmed significant effects across all performance measures.
@@
    • Percentile rank rescaling: Results with percentile rank rescaling of DreamSim scores were consistent with the Z-score findings. DALL-E 3 (Verbatim) participants outperformed both other conditions on first and average performance, with no differences between DALL-E 3 (Revised) and DALL-E 2. For best attempts, all pairwise comparisons were significant, showing DALL-E 3 (Verbatim) ¿ DALL-E 3 (Revised) ¿ DALL-E 2. ANOVA tests confirmed significant effects across all performance measures.
  • Prompting Behaviors
@@
    • Mean prompt Length: We found statistically significant differences in prompt length across treatment arms. Participants using DALL-E 2 wrote significantly shorter prompts compared to both DALL-E 3 (Revised) and DALL-E 3 (Verbatim) groups. Interestingly, there was no significant difference in prompt length between the two DALL-E 3 variants. ANOVA results confirmed a significant effect of model version on prompt length.
@@
    • Aggregate Similarity: We found no statistically significant differences in the overall variability of prompts across the three treatment arms. ANOVA results confirmed no significant effect of model version on this measure, suggesting that participants explored similar breadths of prompt space regardless of the model they were using.
@@
    • Successive Similarity: Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences across treatment arms. ANOVA results confirmed no significant effect of model version on successive prompt similarity, indicating that the tendency to build upon or deviate from previous prompts was similar across models.
@@
    • Successive Prompt Token Sort Ratio: No statistically significant differences emerged across treatment arms. ANOVA results confirmed no significant effect of model version on this measure of textual similarity between successive prompts.
@@
    • Successive Prompt ”Contains Previous Prompt” Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences across treatment arms. ANOVA results confirmed no significant effect of model version on this measure.
@@
Analysis approach: To test this hypothesis, we conducted two types of analyses. First, we performed F-tests comparing the variance of participant performance and prompting behaviors between all three pairs of model conditions (DALL-E 2 vs. DALL-E 3 Revised, DALL-E 3 Revised vs. DALL-E 3 Verbatim, and DALL-E 2 vs. DALL-E 3 Verbatim). Second, we estimated quantile treatment effects (QTEs) between all three pairs of models on participant performance and prompting behaviors. We also visually inspected the QTE patterns to determine whether dispersion/”inequality” was being reduced or increased when participants used different models. For example, positive effects for low quantiles and negative/null effects for high quantiles would indicate inequality reduction.
F-test Results (with BH Adjusted p-values):
  • DALL-E 3 (revised) vs. DALL-E 2
@@
  • DALL-E 3 (verbatim) vs. DALL-E 3 (revised)
@@
  • DALL-E 2 vs. DALL-E 3 (verbatim)
@@
QTE Highlighted Results: Through our quantile treatment effects analysis and visual inspection of QTE plots, we found:
  • Evidence of dispersion reduction: Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALL-E 3 Revised vs. DALL-E 2 comparison showed patterns consistent with inequality reduction, with larger positive effects at lower quantiles.
  • Evidence of dispersion increase: Mean prompt length for all three pairwise model comparisons (DALL-E 3 Verbatim vs. DALL-E 2, DALL-E 3 Revised vs. DALL-E 2, and DALL-E 3 Verbatim vs. DALL-E 3 Revised) showed patterns consistent with increased inequality, with larger positive effects at higher quantiles. Similar patterns emerged for raw DreamSim performance in both DALL-E 3 Verbatim vs. DALL-E 2 and DALL-E 3 Revised vs. DALL-E 2 comparisons.
@@
Analysis approach: To evaluate how performance improves across successive attempts and whether this improvement varies by model, we conducted stratified two-sample tests comparing participants’ initial scores with their best scores. These comparisons were performed both within each treatment arm (DALL-E 2, DALL-E 3 Verbatim, and DALL-E 3 Revised) and overall across all participants. This approach allowed us to determine both whether participants generally improved with practice and whether the rate of improvement differed across models with varying capabilities.
Results: Our analyses revealed consistent performance improvements across all treatment arms and outcome measures:
  • CLIP embedding cosine similarity
    • No rescaling: We found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample. This indicates that participants were able to improve their performance through iterative attempts regardless of which model they were using.
    • Z-score rescaling: Results mirrored those of the unscaled analysis, with statistically significant improvements from initial to best performance observed both within each treatment arm and overall.
    • Percentile rank rescaling: Consistent with other scaling approaches, we observed statistically significant improvements from initial to best performance both within each treatment arm and overall.
  • DreamSim
    • No rescaling: DreamSim results aligned with CLIP similarity findings, showing statistically significant improvements from initial to best performance both within each treatment arm and overall.
    • Z-score rescaling: We again found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample.
    • Percentile rank rescaling: Results were consistent with other analyses, showing statistically significant improvements from initial to best performance across all conditions.
@@
Analysis approach: To evaluate this hypothesis, we employed two complementary approaches. First, we used GPT-4V (a multimodal generative AI model capable of processing both text and images) to generate optimal prompts for each target image. For each of our 15 target images, we instructed GPT-4V to ”Write a DALL-E 2, 3 prompt to recreate this image verbatim as closely and as detailed as possible.” We generated two such ”AI prompts” per target image—one optimized for DALL-E 2 and one for DALL-E 3. We then submitted these AI-generated prompts to the respective models 20 times each, yielding 60 replicated images per target (as the DALL-E 3 prompts were sent to both DALL-E 3 variants). We measured the cosine similarity between CLIP embeddings of these generated images and their target images, then averaged these similarity scores to quantify GPT-4V’s ability to generate effective replication prompts for each target.
@@
Results:
  • CLIP embedding cosine similarity: When ranking target images by GPT-4V’s ability to generate effective prompts, we found substantial variation across images:
@@ Results:
  • DreamSim: When using DreamSim as our similarity metric (inverted as 1-DreamSim to create a similarity rather than distance measure), GPT-4V’s prompt generation performance ranked as follows:
@@
”We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate. We anticipate power for this analysis will be very low, so we chose to label it as an exploratory analysis rather than a pre-registered hypothesis.
@@
G.2 Deviations From Pre-registration
@@
  • Statistical tests: Our pre-registration specified t-tests and Mann-Whitney U tests for many hypotheses. However, this approach proved inappropriate for variables with multiple categories (which characterized most demographic traits). We therefore employed ANOVA and Kruskal-Wallis tests instead, which provide equivalent information for multi-category variables.
  • Z-score computation: Our pre-registration was insufficiently precise about computing Z-score performance measures. As detailed in Section D.2, we calculated Z-scores within image-attempt pairs to adjust for variations across target images and attempts, except when comparing performance across attempts, where Z-scores were computed within target images only. The pre-registration had anticipated computing Z-scores only within target images for all analyses.
  • Additional demographic variables: We inadvertently omitted Education and Generative AI outlook from our pre-registered list of demographics for measuring task performance heterogeneity. Given their theoretical importance, we included these variables in our analyses despite this accidental omission.
  • Additional prompt exclusion criteria: Beyond our pre-registered exclusion criteria, we removed additional prompts that did not appear to constitute ”good-faith efforts” based on their text content (see Section C.4 for details). Robustness checks confirmed that our results remain consistent when including these prompts.
  • T-tests vs. Z-tests: For testing H5, we used t-tests rather than z-tests because they were easier to implement. As these tests are asymptotically equivalent, this change does not meaningfully affect our results.
  • Z-score outlier handling: When testing H4, we observed that Z-scores did not follow a normal distribution and included extreme values (ranging from -21.9 to 7.1). To prevent these outliers from disproportionately influencing our linear models, we excluded observations with absolute Z-scores greater than 3 (2.65
  • GPT-4V rescaling omission: Our pre-registration included a rescaled version of performance metrics based on GPT-4V prompt quality. However, this rescaling would have amounted to subtracting a constant from each unscaled score, and since all our analyses adjusted for target image (either through post-stratification or as a covariate), results would have been identical to those using unscaled measures. We therefore omitted these redundant robustness checks.
G.3 OSF Pre-registration
A complete copy of the pre-registration deposited on OSF on December 11, 2023 will be included beginning on the next page in the final manuscript, but has been removed here to maintain anonymity.

### v6

@@ I.1 Hypotheses and Results
CLIP Z-score rescaling.
@@
We found statistically significant evidence for performance differences across treatment arms when examining participants’ first attempts, average performance across all attempts, and best attempts (all measured using taskwide Z-scores rather than task-iteration Z-scores). For first attempt performance, DALL-E 3 (Verbatim) participants significantly outperformed both DALL-E 3 (Revised) and DALL-E 2 participants, while no statistically significant difference emerged between DALL-E 3 (Revised) and DALL-E 2 participants. Average performance across all attempts showed the same pattern: DALL-E 3 (Verbatim) outperformed both other conditions, with no significant difference between DALL-E 3 (Revised) and DALL-E 2. When comparing participants’ best attempts, we found statistically significant differences between all three treatment arms in a clear hierarchy: DALL-E 3 (Verbatim) produced better results than DALL-E 3 (Revised), which in turn produced better results than DALL-E 2. ANOVA tests for all three performance measures (first, average, and best attempts) confirmed significant effects of model assignment.
CLIP Percentile rank rescaling.
@@
Results using percentile rank rescaling mirrored those found with Z-score rescaling. DALL-E 3 (Verbatim) participants outperformed both other conditions on first attempt performance and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2 for these metrics. For best attempt performance, all pairwise comparisons revealed significant differences, with DALL-E 3 (Verbatim) outperforming DALL-E 3 (Revised), which outperformed DALL-E 2. ANOVA tests confirmed significant effects of model assignment for all three performance measures.
DreamSim Z-score rescaling.
@@
Results using DreamSim with Z-score rescaling showed the same pattern as CLIP embedding cosine similarity. DALL-E 3 (Verbatim) participants significantly outperformed both other conditions on first attempt and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2. For best attempts, we again found the same hierarchical pattern of performance differences across all three conditions. ANOVA tests confirmed significant effects across all performance measures.
DreamSim Percentile rank rescaling.
@@
Results with percentile rank rescaling of DreamSim scores were consistent with the Z-score findings. DALL-E 3 (Verbatim) participants outperformed both other conditions on first and average performance, with no differences between DALL-E 3 (Revised) and DALL-E 2. For best attempts, all pairwise comparisons were significant, showing DALL-E 3 (Verbatim) ¿ DALL-E 3 (Revised) ¿ DALL-E 2. ANOVA tests confirmed significant effects across all performance measures.
Mean prompt Length.
@@
We found statistically significant differences in prompt length across treatment arms. Participants using DALL-E 2 wrote significantly shorter prompts compared to both DALL-E 3 (Revised) and DALL-E 3 (Verbatim) groups. Interestingly, there was no significant difference in prompt length between the two DALL-E 3 variants. ANOVA results confirmed a significant effect of model version on prompt length.
Aggregate Similarity.
@@
We found no statistically significant differences in the overall variability of prompts across the three treatment arms. ANOVA results confirmed no significant effect of model version on this measure, suggesting that participants explored similar breadths of prompt space regardless of the model they were using. These results contrast with our findings from the main analysis, where prompts written by DALL-E 3 participants were, on average, more consistent than those written by DALL-E 2 participants. This difference reflects statistical power and specification: the main analysis uses a higher-powered model with task and attempt fixed effects, whereas the aggregate-similarity analysis here relies on a simpler cross-sectional comparison.
Successive Similarity.
@@
Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences across treatment arms. ANOVA results confirmed no significant effect of model version on successive prompt similarity, indicating that the tendency to build upon or deviate from previous prompts was similar across models.
Successive prompt token sort ratio.
@@
No statistically significant differences emerged across treatment arms. ANOVA results confirmed no significant effect of model version on this measure of textual similarity between successive prompts.
Successive prompt ‘contains previous prompt’ dummy:
@@
Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences across treatment arms. ANOVA results confirmed no significant effect of model version on this measure.
Interpretation.
@@
Empirical approach:
To test this hypothesis, we conducted two types of analyses. First, we performed F-tests comparing the variance of participant performance and prompting behaviors between all three pairs of model conditions (DALL-E 2 vs. DALL-E 3 Revised, DALL-E 3 Revised vs. DALL-E 3 Verbatim, and DALL-E 2 vs. DALL-E 3 Verbatim). Second, we estimated quantile treatment effects (QTEs) between all three pairs of models on participant performance and prompting behaviors. We also visually inspected the QTE patterns to determine whether dispersion/”inequality” was being reduced or increased when participants used different models. For example, positive effects for low quantiles and negative/null effects for high quantiles would indicate inequality reduction.
F-tests.
We first present results (with BH Adjusted p-values) for DALL-E 3 (revised) vs. DALL-E 2:
  1. 1.
@@
We next present results (with BH Adjusted p-values) for DALL-E 3 (verbatim) vs. DALL-E 3 (revised):
  1. 1.
@@
Finally, the F-tests for DALL-E 2 vs. DALL-E 3 (verbatim):
  1. 1.
@@
QTE Highlighted Results.
Through our quantile treatment effects analysis and visual inspection of QTE plots, we found clear evidence of dispersion reduction. The Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALL-E 3 Revised vs. DALL-E 2 comparison showed patterns consistent with inequality reduction, with larger positive effects at lower quantiles. We also found evidence of dispersion increase: Mean prompt length for all three pairwise model comparisons (DALL-E 3 Verbatim vs. DALL-E 2, DALL-E 3 Revised vs. DALL-E 2, and DALL-E 3 Verbatim vs. DALL-E 3 Revised) showed patterns consistent with increased inequality, with larger positive effects at higher quantiles. Similar patterns emerged for raw DreamSim performance in both DALL-E 3 Verbatim vs. DALL-E 2 and DALL-E 3 Revised vs. DALL-E 2 comparisons.
@@
Empirical approach:
To evaluate how performance improves across successive attempts and whether this improvement varies by model, we conducted stratified two-sample tests comparing participants’ initial scores with their best scores. These comparisons were performed both within each treatment arm (DALL-E 2, DALL-E 3 Verbatim, and DALL-E 3 Revised) and overall across all participants. This approach allowed us to determine both whether participants generally improved with practice and whether the rate of improvement differed across models with varying capabilities.
Our analyses revealed consistent performance improvements across all treatment arms and outcome measures.
CLIP No rescaling.
We found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample. This indicates that participants were able to improve their performance through iterative attempts regardless of which model they were using.
CLIP Z-score rescaling.
Results mirrored those of the unscaled analysis, with statistically significant improvements from initial to best performance observed both within each treatment arm and overall.
CLIP Percentile rank rescaling.
Consistent with other scaling approaches, we observed statistically significant improvements from initial to best performance both within each treatment arm and overall.
DreamSim No rescaling.
DreamSim results aligned with CLIP similarity findings, showing statistically significant improvements from initial to best performance both within each treatment arm and overall.
DreamSim Z-score rescaling.
We again found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample.
DreamSim Percentile rank rescaling.
Results were consistent with other analyses, showing statistically significant improvements from initial to best performance across all conditions.
Interpretation.
@@
Empirical approach.
To evaluate this hypothesis, we employed two complementary approaches. First, we used GPT-4V (a multimodal generative AI model capable of processing both text and images) to generate optimal prompts for each target image. For each of our 15 target images, we instructed GPT-4V to ”Write a DALL-E 2, 3 prompt to recreate this image verbatim as closely and as detailed as possible.” We generated two such ”AI prompts” per target image—one optimized for DALL-E 2 and one for DALL-E 3. We then submitted these AI-generated prompts to the respective models 20 times each, yielding 60 replicated images per target (as the DALL-E 3 prompts were sent to both DALL-E 3 variants). We measured the cosine similarity between CLIP embeddings of these generated images and their target images, then averaged these similarity scores to quantify GPT-4V’s ability to generate effective replication prompts for each target.
@@
CLIP Results.
When ranking target images by GPT-4V’s ability to generate effective prompts, we found substantial variation across images:
@@
DreamSim Results.
When using DreamSim as our similarity metric (inverted as 1-DreamSim to create a similarity rather than distance measure), GPT-4V’s prompt generation performance ranked as follows:
@@
“We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate. We anticipate power for this analysis will be very low, so we chose to label it as an exploratory analysis rather than a pre-registered hypothesis.
@@
I.2 Deviations From Pre-registration
@@
Statistical tests.
Our pre-registration specified t-tests and Mann-Whitney U tests for many hypotheses. However, this approach proved inappropriate for variables with multiple categories (which characterized most demographic traits). We therefore employed ANOVA and Kruskal-Wallis tests instead, which provide equivalent information for multi-category variables.
Z-score computation.
Our pre-registration was insufficiently precise about computing Z-score performance measures. As detailed in Section E.2, we calculated Z-scores within image-attempt pairs to adjust for variations across target images and attempts, except when comparing performance across attempts, where Z-scores were computed within target images only. The pre-registration had anticipated computing Z-scores only within target images for all analyses.
Additional demographic variables.
We inadvertently omitted Education and Generative AI outlook from our pre-registered list of demographics for measuring task performance heterogeneity. Given their theoretical importance, we included these variables in our analyses despite this accidental omission.
Additional prompt exclusion criteria.
Beyond our pre-registered exclusion criteria, we removed additional prompts that did not appear to constitute ”good-faith efforts” based on their text content (see Section C.4 for details). Robustness checks confirmed that our results remain consistent when including these prompts.
T-tests vs. Z-tests.
For testing H5, we used t-tests rather than z-tests because they were easier to implement. As these tests are asymptotically equivalent, this change does not meaningfully affect our results.
Z-score outlier handling.
When testing H4, we observed that Z-scores did not follow a normal distribution and included extreme values (ranging from -21.9 to 7.1). To prevent these outliers from disproportionately influencing our linear models, we excluded observations with absolute Z-scores greater than 3 (2.65% of user-attempt observations). Robustness checks including these outliers yielded nearly identical results, with minor differences in significance levels for some comparisons.
GPT-4V rescaling omission.
Our pre-registration included a rescaled version of performance metrics based on GPT-4V prompt quality. However, this rescaling would have amounted to subtracting a constant from each unscaled score, and since all our analyses adjusted for target image (either through post-stratification or as a covariate), results would have been identical to those using unscaled measures. We therefore omitted these redundant robustness checks.